# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ 1 Introduction
Quantum teleportation is one of the most striking applications of quantum entanglement, enabling the transfer of an unknown quantum state from one system to another using entanglement and feedforward operations [1, 2]. It is a fundamental operation that serves as a cornerstone for numerous quantum protocols, underpinning advances in fields such as quantum communication and quantum computing. It plays a crucial role in enabling secure information transfer, distributed quantum networks, and the construction of scalable quantum architectures by facilitating state transfer without the need for direct physical transport of qubits.
@@
Since its initial experimental demonstration with optical photons in Ref. [3], a wide range of quantum teleportation protocols have been proposed and implemented. The most basic version, involving qubit-based systems, has since been realized on numerous platforms [4, 5, 6, 7]. The process relies on entangled pairs of qubits being utilized as a resource for transferring quantum states. In Ref. [8], the authors demonstrated fault-tolerant quantum state teleportation on a trapped-ion platform, marking significant progress toward universal quantum computing. In Ref. [9], the authors demonstrated a teleportation fidelity of over 90%, using time-bin encoded qubits. Recently, there has been increasing interest in extending teleportation to high-dimensional systems [10, 11, 12, 13, 14] and encoded systems [15, 16, 17, 18].
@@
In parallel to teleportation in discrete-variable systems, protocols for continuous-variable (CV) systems were also theoretically and experimentally developed [19, 20, 21, 22, 23]. Here, the scheme typically involves sharing a CV-entangled state, such as a two-mode squeezed state, and quadrature measurements via homodyne detection are performed to transmit quantum information. CV quantum teleportation was experimentally demonstrated for the first time by Furusawa and colleagues [24], realizing unconditional teleportation. Subsequently, several variants of the CV teleportation were explored. In Ref. [25], the authors proposed a hybrid approach combining discrete and continuous variable techniques. Moreover, CV quantum teleportation between atomic ensembles has been demonstrated [26], and the teleportation of four degrees of freedom in a system of optical modes in Ref. [23]. These efforts motivate the exploration of quantum teleportation protocols in other systems.
In standard CV teleportation, quadrature measurements are a crucial step to transmit the unknown state from Alice to Bob. While this can be practically implemented experimentally in optical systems, for example, they are not the natural type of measurements performed in other systems such as microwave cavity modes in circuit quantum electrodynamics (QED), vibration modes in circuit quantum acoustodynamics (QAD), or motional states of trapped ions. These systems are described by the Jaynes-Cummings (JC) Hamiltonian [27, 28, 29], where it is straightforward to perform binary-outcome measurements on the bosonic mode through a two-level system (qubit) to which it couples. For this reason, it is highly desirable to have a teleportation protocol which uses the state preparation and measurement toolbox enabled by the JC interaction. This would also have direct applications to CV quantum information processing, where logical qubits encoded on cat states with different parity can be transferred from one quantum system to another.
@@
This paper is organized as follows. In Sec. 2, we introduce the experimental system and operations that we assume are available to perform the teleportation protocol. In Sec. 3, we describe and analyze the entire quantum teleportation protocol analytically. In Sec. 4, we show numerical simulation results based on the single parity measurement version of the teleportation protocol. In Sec. 5 we show the corresponding simulation results for joint parity measurements. Finally, in Sec. 6 we summarize and discuss our results.
2 Experimental system and available operations
The experimental platform under consideration consists of a physical system with three bosonic modes and at least one qubit, interacting via a Jaynes-Cummings Hamiltonian (Fig. 1). Concrete examples of the physical system include circuit-QED [27] or QAD [28, 30] where different microwave cavity modes or acoustic vibration modes are coupled to a superconducting qubit. Alternatively, trapped ions [29] where motional degrees of freedom are coupled to internal (spin) states are another potential realization. In these systems, the oscillatory modes are treated as independent and are typically distinguished by their frequencies.
@@ 2.1 Preparation of states
In our work, we will consider the teleportation in the basis of Schrödinger cat states, namely coherent superpositions of coherent states that take the form |$\ket{\alpha}+e^{i\phi}\ket{-\alpha}$. These can be experimentally prepared using different approaches, such as through conditional-displacement gates [31, 32], parity measurements [33], or nonlinear evolutions [34, 35, 36]. As an example of the latter approach, it is possible to generate a cat state by interacting the mode with a qubit that follows the JC interaction
@@
for a time $t^{\ast}=\pi\alpha/g$, where the initial state is $|0\rangle|\alpha\rangle$ and $g$ is the coupling strength. Here $\sigma^{\pm}=(\sigma^{x}\pm i\sigma^{y})/2$ and $\sigma^{x},\sigma^{y},\sigma^{z}$ are Pauli matrices. At time $t^{\ast}$ the qubit and mode disentangle, leaving the mode in (approximately) a Schrödinger cat state [37]
@@
2.2 Mode transformations
Teleportation protocols require operations between bosonic modes, which can also be implemented through a (multimode) JC Hamiltonian. In particular, the nonlinearity of an off-resonantly coupled qubit can be used to mediate interactions between different modes through parametric interactions. For example, effective beam-splitter (BS) transformations have been demonstrated between motional states of trapped ions [38], microwave modes in cQED [39] and acoustic vibration modes in cQAD [30]. The resulting BS mode transformations are given by
@@ 2.2 Mode transformations
2.3 Measurement of modes
By working in the strong dispersive regime of the JC Hamiltonian, it is possible to perform parity measurements of the mode [40, 32, 28]. The effective Hamiltonian for the single mode $f$ in the dispersive regime reads
@@ 2.3 Measurement of modes
where the phonon frequency $\omega$ is set to $\frac{|g^{2}|}{\Delta}$, and $G$ is the dispersive coupling parameter $G=\frac{|g|^{2}}{\Delta}$. Using the qubit as the measurement pointer, this Hamiltonian allows us to perform a measurement of the mode. Initializing the qubit in the state $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$, interacting the Hamiltonian for a time $t$, and projecting the qubit in the $\sigma^{x}$ eigenbasis leaves the modes in a state that is equivalent to applying the measurement operators [41]
@@
3 Teleportation protocol
We now describe our quantum teleportation protocol. The overall scheme is summarized schematically in Fig. 2. We draw the protocol in the style of a optical circuit, but it should be understood that this is a schematic representation, and does not constitute how the experiment physically appears. Alice is in possession of the two modes $a,b$ and Bob has the mode $c$, and the aim is to teleport the state initially on mode $a$ to mode $c$. In a similar way to CV teleportation [42], the procedure consists of two beam splitter operations followed by a sequence of measurements and conditional operations [20]. As with the CV case, the role of the first beam splitter is to create a shared entangled state between Alice and Bob. This is achieved by applying a beam splitter to a cat state. The measurements consist of parity measurements on the output modes of the second beam splitter, and a further dispersive measurement is made following the parity measurement. Classically transmitting the results of the measurements to Bob, correction operations can be applied to complete the teleportation operation. We now describe step by step how this procedure achieves quantum teleportation.
3.1 State Initialization
@@
where |±$|\pm\alpha\rangle$ are coherent states. Throughout this paper we consider $\alpha$ to be a real parameter ($\alpha\in\mathbb{R}$). Eq. (10) are even and odd cat states as they respectively consist only of Fock states with either even or odd parity. As such they satisfy the orthogonality relation
@@
|±$|\pm\alpha\rangle$ $=\sqrt{\frac{(1+e^{-2|\alpha|^{2}})}{2}}|0_{L}\rangle_{\alpha}\pm\sqrt{\frac{(1-e^{-2|\alpha|^{2}})}{2}}|1_{L}\rangle_{\alpha}.$ (13)
@@
|±$|\pm\alpha\rangle$ $\approx\frac{1}{\sqrt{2}}(|0_{L}\rangle_{\alpha}\pm|1_{L}\rangle_{\alpha}).$ (14)
@@
3.2 Generation of a shared entangled state
@@
where we used (14). We see that an entangled state is created between Alice and Bob’s modes. Superscripts in Eq. (19) indicate the mode for the state. Note that this state can also be prepared through conditional displacement operations [43].
3.3 Second beam splitter
@@
The transformed state (20) is reminiscent of the state prior to measurement in standard qubit teleportation [44]. The state consists of a Bell state on Alice’s modes and variations of the initial state on Bob’s side. A measurement of the first two modes collapses the state to one of the four terms in the superposition, leaving the remaining mode in one of four states that correspond to a bit or phase flipped version of Alice’s original state. The task is now to perform a measurement of the first two modes in such a way to induce this collapse using only parity and dispersive measurements.
3.4 Measurement
@@
3.4.1 Variant 1: Single-mode parity measurements
@@
Figure 3: The various coherent states that appear in our quantum teleportation protocol. (a)(c) The coherent states |±$|\pm\alpha\rangle$ form the original logical states on mode $a$, where we take $\alpha$ to be real. The coherent states |±$|\pm\beta\rangle$ form the Bell state on modes $b$ and $c$ and can be taken to be either (a) purely imaginary or (c) purely real. The combinations of these |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ as defined in (21) resulting from the second beam splitter is also shown in each case. (b)(d) The amplitudes of the |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ states after performing a displacement $D(\chi)$. In (b), the inner dashed line shows the amplitude of the states $|\bar{\chi}|=\sqrt{2}\alpha$. The outer dashed line is the amplitude of the state $|\chi\rangle$ which has a radius $2\alpha$. (e)(f)The probability distributions in the Fock basis $p_{n}=|\langle n|\xi\rangle|^{2}$ for the coherent states with $\xi=0,2\chi,\chi\pm\bar{\chi}$. Parameters chosen are (e) $\alpha=5,\beta=5i$; (f) $\alpha=5,\beta=5$.
@@
A plot of the coherent states that appear in our protocol are shown in Fig. 3. We see that for the choice of $\alpha>0$ real and purely imaginary $\beta=i\alpha$, the coherent states |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ lie in distinct parts of phase space to each other (Fig. 3(a)). Alternatively, for the choice of $\alpha\sim\beta$ both real, |±$|\pm\chi\rangle$ appear in distinct parts of phase space, but |±$|\pm\bar{\chi}\rangle$ tends to be close together. If we are able to perform a measurement identifying whether mode $a$ has amplitude ±$\pm\chi$ or ±$\pm\bar{\chi}$, we should be able to collapse the state (26) further and obtain the teleported state.
@@
Such a measurement can be accomplished by performing a displacement $D(\chi)$ followed by a measurement in the Fock basis. Fig. 3(b)(d) shows the amplitude of the |±$|\pm\chi\rangle$ or |±$|\pm\bar{\chi}\rangle$ states after the displacement. For the choice $\beta=i\alpha$ (Fig. 3(b)), we see that the |±$|\pm\chi\rangle$ states now have amplitudes of 00 and $2\alpha$ respectively, while the |±$|\pm\bar{\chi}\rangle$ states both have an amplitude of $\sqrt{2}\alpha$. As these have different amplitudes, an estimate of the Fock number will collapse the state to one of the coherent states. For $\beta=\alpha$ taking real values, the |±$|\pm\chi\rangle$ states now have amplitudes of 00 and $2\sqrt{2}\alpha$, while the |±$|\pm\bar{\chi}\rangle$ states have an amplitude of $\sqrt{2}\alpha$.
@@
The number distribution of the states |±$|\pm\chi\rangle,|\pm\bar{\chi}\rangle$ after a displacement of $D(\chi)$ is shown in Figs. 3(e)(f). For the $\beta=i\alpha$ case (Fig. 3(e)), we see that there is a sharp peak at $n=0$ for the vacuum state, as well as peaks at $n=2\alpha^{2}$ and $n=4\alpha^{2}$ for the states |±$|\pm\bar{\chi}\rangle$ and $|+\chi\rangle$ respectively. Therefore, taking as an example the state with $\sigma_{a}=\sigma_{b}=+$ in Eq. (26), measuring a Fock state $|n\rangle$ with $n\approx 2\alpha^{2}$ will collapse the state in modes $c$ to $\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$, while $n=0$ or $n\approx 4\alpha^{2}$ will collapse the state in modes $c$ to $\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$. For the $\beta=\alpha$ case, we see that there is a sharp peak at $n=0$ for the vacuum state (Fig. 3(f)), as well as peaks at $n=2\alpha^{2}$ and $n=8\alpha^{2}$ for the states |±$|\pm\bar{\chi}\rangle$ and $|+\chi\rangle$ respectively. Again, a measurement in the Fock basis will distinguish the coherent state amplitudes in this case.
@@
$\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$ $+$ $+$ $+$ $I_{L}$
$\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$ $+$ $+$ -- $Z_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$ $+$ -- $+$ $X_{L}$
$-\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$ $+$ -- -- $X_{L}Z_{L}$
@@
is the average photon number of the states $|\chi\pm\bar{\chi}\rangle$.
For example, with the choice $\beta=i\alpha$, the optimal time is $\tau=\pi/(4\alpha^{2})$. This choice puts the zero of the cosine function for the $M_{+}$ outcome coinciding with the |±$|\pm\bar{\chi}\rangle$ states (Fig. 3(e)). Meanwhile, the sine function is a zero at location of the |±$|\pm\chi\rangle$ states. Hence, we approximately have for $\beta=i\alpha$:
@@
Meanwhile, for $\beta=\alpha$, the optimal time is again $\tau=\pi/(4\alpha^{2})$. As can be seen from Fig. 3(f), the cosine function again has a zero at the location of the |±$|\pm\bar{\chi}\rangle$ states. The sine function again takes zero values at the location of the |±$|\pm\chi\rangle$ peaks. In this case, we approximately have for $\beta=\alpha$:
@@
Hence, for both $\beta=i\alpha,\alpha$ the displaced measurements are effective at distinguishing between the |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ states.
@@
where $k_{\pm}$ are the number of times the outcome $M_{\pm}$ is obtained, respectively, and the total number of measurements is $N=k_{+}+k_{-}$. The binomial factor is present because there can be more than one outcome that have the same $k_{\pm}$. For example, for $N=2$ measurements, $M_{+}M_{-}$ and $M_{-}M_{+}$ are both outcomes with $k_{+}=k_{-}=1$. The binomial factor ensures that $\sum_{k_{+}=0}^{N}M_{k_{+},N-k_{+}}^{\dagger}M_{k_{+},N-k_{+}}=I$. The effect of multiple measurements is to cause a further collapse on the Fock basis. The trigonometric factor in (32) can be approximated by a Gaussian [45, 46]
($\sqrt{{N\choose k_{+}}}\text{cos}^{k_{+}}(x)\text{sin}^{k_{-}}(x)\propto\exp(-N(x-x_{k_{+}k_{-}})^{2}),$ (33)
@@
Then one can find which Fock state the collapse occured at by the relation $n=x_{k_{+}k_{-}}/\tau$. From this information, one may determine which of the coherent states |±$|\pm\chi\rangle,|\pm\bar{\chi}\rangle$ the collapse occurred at.
3.4.2 Variant 2: Joint parity measurements

### v2

# Quantum teleportation of cat states with binary-outcome measurements

@@ I Introduction
Quantum teleportation is one of the most striking applications of quantum entanglement, enabling the transfer of an unknown quantum state from one system to another using entanglement and feedforward operations Bennett et al. (1993); Karlsson and Bourennane (1998). It is a fundamental operation that serves as a cornerstone for numerous quantum protocols, underpinning advances in fields such as quantum communication and quantum computing. It plays a crucial role in enabling secure information transfer, distributed quantum networks, and the construction of scalable quantum architectures by facilitating state transfer without the need for direct physical transport of qubits.
@@
Since its initial experimental demonstration with optical photons in Ref. Bouwmeester et al. (1997), a wide range of quantum teleportation protocols have been proposed and implemented. The most basic version, involving qubit-based systems, has since been realized on numerous platforms Boschi et al. (1998); Barrett et al. (2004); Jin et al. (2010); Nölleke et al. (2013). The process relies on entangled pairs of qubits being utilized as a resource for transferring quantum states. In Ref. Ryan-Anderson et al. (2024a), the authors demonstrated fault-tolerant quantum state teleportation on a trapped-ion platform, marking significant progress toward universal quantum computing. In Ref. Valivarthi et al. (2020), the authors demonstrated a teleportation fidelity of over 90%, using time-bin encoded qubits. Recently, there has been increasing interest in extending teleportation to high-dimensional systems Luo et al. (2019); Hu et al. (2020); Zhang et al. (2019, 2006); Erhard et al. (2020) and encoded systems Huang et al. (2021); Luo et al. (2021); Chaudhary et al. (2024); Ryan-Anderson et al. (2024b).
@@
In parallel to teleportation in discrete-variable systems, protocols for continuous-variable (CV) systems were also theoretically and experimentally developed Vaidman (1994); Braunstein and Kimble (1998); Bowen et al. (2003); Yukawa et al. (2008); Liu et al. (2024). Here, the scheme typically involves sharing a CV-entangled state, such as a two-mode squeezed state, and quadrature measurements via homodyne detection are performed to transmit quantum information. CV quantum teleportation was experimentally demonstrated for the first time by Furusawa and colleagues Furusawa et al. (1998), realizing unconditional teleportation. Subsequently, several variants of the CV teleportation were explored. In Ref. Lee et al. (2011), the authors proposed a hybrid approach combining discrete and continuous variable techniques. Moreover, CV quantum teleportation between atomic ensembles has been demonstrated Krauter et al. (2013), and the teleportation of four degrees of freedom in a system of optical modes in Ref. Liu et al. (2024). These efforts motivate the exploration of quantum teleportation protocols in other systems.
In standard CV teleportation, quadrature measurements are a crucial step to transmit the unknown state from Alice to Bob. While this can be practically implemented experimentally in optical systems, for example, they are not the natural type of measurements performed in other systems such as microwave cavity modes in circuit quantum electrodynamics (QED), vibration modes in circuit quantum acoustodynamics (QAD), or motional states of trapped ions. In fact, in the above systems the microwave modes or motional states are measured by coupling them to a two-level system through a Jaynes-Cummings (JC) interaction Blais et al. (2021); von Lüpke et al. (2022); Leibfried et al. (2003), resulting in the possibility of extracting only one bit of information per measurement repetition. For this reason, it is much easier and time-efficient to extract binary observables such as parity, rather than performing continuous variable measurements of phase-space quadratures. While alternative approaches to measure quadratures e.g. cQED exist, for example through parametrically-engineered longitudinal coupling or optimal control pulses Krisnanda et al. (2025), parity measurements are still more robust and efficient due to their immediate implementation in the dispersive regime of the Jaynes-Cummings Hamiltonian. For this reason, having available a teleportation protocol which relies only on parity measurements is of central interest.
@@
This paper is organized as follows. In Sec. II, we introduce the experimental system and operations that we assume are available to perform the teleportation protocol. In Sec. III, we describe and analyze the entire quantum teleportation protocol analytically. In Sec. IV, we show numerical simulation results based on the single parity measurement version of the teleportation protocol. In Sec. V we show the corresponding simulation results for joint parity measurements. Finally, in Sec. VII we summarize and discuss our results.
II Experimental system and available operations
The experimental platform under consideration consists of a physical system with three bosonic modes and at least one qubit, interacting via a Jaynes-Cummings Hamiltonian (Fig. 1). Concrete examples of the physical system include circuit-QED Blais et al. (2021) or QAD von Lüpke et al. (2022, 2024) where different microwave cavity modes or acoustic vibration modes are coupled to a superconducting qubit. Alternatively, trapped ions Leibfried et al. (2003) where motional degrees of freedom are coupled to internal (spin) states are another potential realization. In these systems, the oscillatory modes are treated as independent and are typically distinguished by their frequencies.
@@ II.1 Preparation of states
[…]
@@
for a time $t^{\ast}=\pi\alpha/g$, where the initial state is $|0\rangle|\alpha\rangle$ and $g$ is the coupling strength. Here $\sigma^{\pm}=(\sigma^{x}\pm i\sigma^{y})/2$ and $\sigma^{x},\sigma^{y},\sigma^{z}$ are Pauli matrices. At time $t^{\ast}$ the qubit and mode disentangle, leaving the mode in (approximately) a Schrödinger cat state Gea-Banacloche (1991)
@@
II.2 Mode transformations
Teleportation protocols require operations between bosonic modes, which can also be implemented through a (multimode) JC Hamiltonian. In particular, the nonlinearity of an off-resonantly coupled qubit can be used to mediate interactions between different modes through parametric interactions. For example, effective beam-splitter (BS) transformations have been demonstrated between motional states of trapped ions Leibfried et al. (2002), microwave modes in cQED Gao et al. (2018) and acoustic vibration modes in cQAD von Lüpke et al. (2024). The resulting BS mode transformations are given by
@@ II.2 Mode transformations
II.3 Measurement of modes
By working in the strong dispersive regime of the JC Hamiltonian, it is possible to perform parity measurements of the mode Lutterbach and Davidovich (1997); Vlastakis et al. (2013); von Lüpke et al. (2022). The effective Hamiltonian for the single mode $f$ in the dispersive regime reads
@@ II.3 Measurement of modes
where the phonon frequency $\omega$ is set to $\frac{|g^{2}|}{\Delta}$, and $G$ is the dispersive coupling parameter $G=\frac{|g|^{2}}{\Delta}$. Using the qubit as the measurement pointer, this Hamiltonian allows us to perform a measurement of the mode. Initializing the qubit in the state $|+\rangle=(|0\rangle+|1\rangle)/\sqrt{2}$, interacting the Hamiltonian for a time $t$, and projecting the qubit in the $\sigma^{x}$ eigenbasis leaves the modes in a state that is equivalent to applying the measurement operators Mao et al. (2023)
@@
III Teleportation protocol
We now describe our quantum teleportation protocol. The overall scheme is summarized schematically in Fig. 2. We draw the protocol in the style of a optical circuit, but it should be understood that this is a schematic representation, and does not constitute how the experiment physically appears. Alice is in possession of the two modes $a,b$ and Bob has the mode $c$, and the aim is to teleport the state initially on mode $a$ to mode $c$. In a similar way to CV teleportation Braunstein and van Loock (2005), the procedure consists of two beam splitter operations followed by a sequence of measurements and conditional operations Braunstein and Kimble (1998). As with the CV case, the role of the first beam splitter is to create a shared entangled state between Alice and Bob. This is achieved by applying a beam splitter to a cat state. The measurements consist of parity measurements on the output modes of the second beam splitter, and a further dispersive measurement is made following the parity measurement. Classically transmitting the results of the measurements to Bob, correction operations can be applied to complete the teleportation operation. We now describe step by step how this procedure achieves quantum teleportation.
III.1 State Initialization
@@
where $|\pm\alpha\rangle$ are coherent states. Throughout this paper we consider $\alpha$ to be a real parameter ($\alpha\in\mathbb{R}$). Eq. (10) are even and odd cat states as they respectively consist only of Fock states with either even or odd parity. As such they satisfy the orthogonality relation
@@
$|\pm\alpha\rangle$ $=\sqrt{\frac{(1+e^{-2|\alpha|^{2}})}{2}}|0_{L}\rangle_{\alpha}\pm\sqrt{\frac{(1-e^{-2|\alpha|^{2}})}{2}}|1_{L}\rangle_{\alpha}.$ (13)
@@
$|\pm\alpha\rangle$ $\approx\frac{1}{\sqrt{2}}(|0_{L}\rangle_{\alpha}\pm|1_{L}\rangle_{\alpha}).$ (14)
@@
III.2 Generation of a shared entangled state
@@
where we used (14). We see that an entangled state is created between Alice and Bob’s modes. Superscripts in Eq. (19) indicate the mode for the state. Note that this state can also be prepared through conditional displacement operations Wang et al. (2016).
III.3 Second beam splitter
@@
The transformed state (20) is reminiscent of the state prior to measurement in standard qubit teleportation Nielsen and Chuang (2010). The state consists of a Bell state on Alice’s modes and variations of the initial state on Bob’s side. A measurement of the first two modes collapses the state to one of the four terms in the superposition, leaving the remaining mode in one of four states that correspond to a bit or phase flipped version of Alice’s original state. The task is now to perform a measurement of the first two modes in such a way to induce this collapse using only parity and dispersive measurements.
III.4 Measurement
@@
III.4.1 Variant 1: Single-mode parity measurements
@@
Figure 3: The various coherent states that appear in our quantum teleportation protocol. (a)(c) The coherent states $|\pm\alpha\rangle$ form the original logical states on mode $a$, where we take $\alpha$ to be real. The coherent states $|\pm\beta\rangle$ form the Bell state on modes $b$ and $c$ and can be taken to be either (a) purely imaginary or (c) purely real. The combinations of these $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ as defined in (21) resulting from the second beam splitter is also shown in each case. (b)(d) The amplitudes of the $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ states after performing a displacement $D(\chi)$. In (b), the inner dashed line shows the amplitude of the states $|\bar{\chi}|=\sqrt{2}\alpha$. The outer dashed line is the amplitude of the state $|\chi\rangle$ which has a radius $2\alpha$. (e)(f)The probability distributions in the Fock basis $p_{n}=|\langle n|\xi\rangle|^{2}$ for the coherent states with $\xi=0,2\chi,\chi\pm\bar{\chi}$. Parameters chosen are (e) $\alpha=5,\beta=5i$; (f) $\alpha=5,\beta=5$.
@@
A plot of the coherent states that appear in our protocol are shown in Fig. 3. We see that for the choice of $\alpha>0$ real and purely imaginary $\beta=i\alpha$, the coherent states $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ lie in distinct parts of phase space to each other (Fig. 3(a)). Alternatively, for the choice of $\alpha\sim\beta$ both real, $|\pm\chi\rangle$ appear in distinct parts of phase space, but $|\pm\bar{\chi}\rangle$ tends to be close together. If we are able to perform a measurement identifying whether mode $a$ has amplitude $\pm\chi$ or $\pm\bar{\chi}$, we should be able to collapse the state (26) further and obtain the teleported state.
@@
Such a measurement can be accomplished by performing a displacement $D(\chi)$ followed by a measurement in the Fock basis. Fig. 3(b)(d) shows the amplitude of the $|\pm\chi\rangle$ or $|\pm\bar{\chi}\rangle$ states after the displacement. For the choice $\beta=i\alpha$ (Fig. 3(b)), we see that the $|\pm\chi\rangle$ states now have amplitudes of 0 and $2\alpha$ respectively, while the $|\pm\bar{\chi}\rangle$ states both have an amplitude of $\sqrt{2}\alpha$. As these have different amplitudes, an estimate of the Fock number will collapse the state to one of the coherent states. For $\beta=\alpha$ taking real values, the $|\pm\chi\rangle$ states now have amplitudes of 0 and $2\sqrt{2}\alpha$, while the $|\pm\bar{\chi}\rangle$ states have an amplitude of $\sqrt{2}\alpha$.
@@
The number distribution of the states $|\pm\chi\rangle,|\pm\bar{\chi}\rangle$ after a displacement of $D(\chi)$ is shown in Figs. 3(e)(f). For the $\beta=i\alpha$ case (Fig. 3(e)), we see that there is a sharp peak at $n=0$ for the vacuum state, as well as peaks at $n=2\alpha^{2}$ and $n=4\alpha^{2}$ for the states $|\pm\bar{\chi}\rangle$ and $|+\chi\rangle$ respectively. Therefore, taking as an example the state with $\sigma_{a}=\sigma_{b}=+$ in Eq. (26), measuring a Fock state $|n\rangle$ with $n\approx 2\alpha^{2}$ will collapse the state in modes $c$ to $\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$, while $n=0$ or $n\approx 4\alpha^{2}$ will collapse the state in modes $c$ to $\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$. For the $\beta=\alpha$ case, we see that there is a sharp peak at $n=0$ for the vacuum state (Fig. 3(f)), as well as peaks at $n=2\alpha^{2}$ and $n=8\alpha^{2}$ for the states $|\pm\bar{\chi}\rangle$ and $|+\chi\rangle$ respectively. Again, a measurement in the Fock basis will distinguish the coherent state amplitudes in this case.
@@
$\mu|0_{L}\rangle_{\beta}+\nu|1_{L}\rangle_{\beta}$ ++ ++ ++ $I_{L}$
$\mu|0_{L}\rangle_{\beta}-\nu|1_{L}\rangle_{\beta}$ ++ ++ - $Z_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$ ++ - ++ $X_{L}$
$\mu|1_{L}\rangle_{\beta}+\nu|0_{L}\rangle_{\beta}$ ++ - - $X_{L}Z_{L}$
@@
is the average boson number of the states $|\chi\pm\bar{\chi}\rangle$.
For example, with the choice $\beta=i\alpha$, the optimal time is $\tau=\pi/(4\alpha^{2})$. This choice puts the zero of the cosine function for the $M_{+}$ outcome coinciding with the $|\pm\bar{\chi}\rangle$ states (Fig. 3(e)). Meanwhile, the sine function is a zero at location of the $|\pm\chi\rangle$ states. Hence, we approximately have for $\beta=i\alpha$:
@@
Meanwhile, for $\beta=\alpha$, the optimal time is again $\tau=\pi/(4\alpha^{2})$. As can be seen from Fig. 3(f), the cosine function again has a zero at the location of the $|\pm\bar{\chi}\rangle$ states. The sine function again takes zero values at the location of the $|\pm\chi\rangle$ peaks. In this case, we approximately have for $\beta=\alpha$:
@@
Hence, for both $\beta=i\alpha,\alpha$ the displaced measurements are effective at distinguishing between the $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ states.
@@
where $k_{\pm}$ are the number of times the outcome $M_{\pm}$ is obtained, respectively, and the total number of measurements is $N=k_{+}+k_{-}$. The binomial factor is present because there can be more than one outcome that have the same $k_{\pm}$. For example, for $N=2$ measurements, $M_{+}M_{-}$ and $M_{-}M_{+}$ are both outcomes with $k_{+}=k_{-}=1$. The binomial factor ensures that $\sum_{k_{+}=0}^{N}M_{k_{+},N-k_{+}}^{\dagger}M_{k_{+},N-k_{+}}=I$. The effect of multiple measurements is to cause a further collapse on the Fock basis. The trigonometric factor in (32) can be approximated by a Gaussian Ilo-Okeke and Byrnes (2016); Kondappan et al. (2023)
$\sqrt{{N\choose k_{+}}}\text{cos}^{k_{+}}(x)\text{sin}^{k_{-}}(x)\propto\exp(-N(x-x_{k_{+}k_{-}})^{2}),$ (33)
@@
Then one can find which Fock state the collapse occured at by the relation $n=x_{k_{+}k_{-}}/\tau$. From this information, one may determine which of the coherent states $|\pm\chi\rangle,|\pm\bar{\chi}\rangle$ the collapse occurred at.
III.4.2 Variant 2: Joint parity measurements